OpenStudy (anonymous):

How to show that (kn)! is divisible by (k!)^n for all k,n >= 1?

OpenStudy (anonymous):

I tried induction but I got nowhere XD

OpenStudy (anonymous):

I have an idea. It is true that product of n consecutive integers is divisible by n!. Then (kn)! is k products of n consecutive integers so (kn)! is divisible by (k!)^n.

OpenStudy (anonymous):

For example for n = 3 and k = 5 I get (5!)^3 = (1 * 2 * 3 * 4 * 5)*(1*2*3*4*5)*(1*2*3*4*5). (5*3)! = (15* 14 * 13 * 12 * 11)*(10*9*8*7*6)*(5*4*3*2*1). then each group of 5 consecutive integers is divisible by 5! and there are k such grups so it is divisible by (5!)^3. Am I right?

OpenStudy (ybarrap):

$$ (kn)!=kn\times(kn-1)\times(kn-2)\cdot\cdot\cdot\times1\\ (k!)^n=\left (k\times(k-1)\times(k-2)\cdot\cdot\cdot\times1\right )^n\\ $$ so $$ \cfrac{(kn)!}{(k)!^n}\\ =\cfrac{kn\times(kn-1)\times(kn-2)\cdot\cdot\cdot\times1}{\left (k\times(k-1)\times(k-2)\cdot\cdot\cdot\times1\right )^n}\\ $$ Let n be arbitray and let \(k\to\infty \): $$ \lim_{k\to\infty}\cfrac{(kn)!}{(k!)^n}=\cfrac{(kn)^n}{k^n}=n^n<\infty\\ $$ Therefore, for any k,n>1, \((kn)!\) is divisible by \((k!)^n\)l